# SEEA-R1: Tree-Structured Reinforcement Fine-Tuning for Self-Evolving Embodied Agents

### v1

@@ 4.2.2 Comparative Study of Training Algorithms
4.2.3 Ablation Study: Self-Evolution with MGRM
We conducted an ablation study to demonstrate that our SEEA-R1 can achieve sustained performance improvement by leveraging its internally trained Multi-Modal Generative Reward Model (MGRM), thereby reducing reliance on ground truth (GT)rewards from the simulator. We compared three critical configurations with results shown in Figure 5:
Figure 4: Performance comparison of SEEA-R1 using different optimization algorithms on ALFWorld over training iterations, more detailed figures are provided in Appendix.
  1. 1.
    SEEA-R1 with GT Rewards: Policy optimization is guided directly by GT environment rewards. As expected, this (typically the blue line) yields the highest success rate, establishing an upper bound due to the perfectly accurate reward signal.
  2. 2.
    SEEA-R1 with frozen MGRM: An untrained and frozen MGRM provides reward signals for policy optimization. Notably, the orange line shows a high initial success rate in the first iteration despite the inaccurate reward signal. It’s possible that the initial exploration coincidentally aligned with the flawed reward. Nevertheless, the fundamental unreliability of the frozen MGRM ultimately misguides the learning process, leading to performance degradation in later iterations.
  3. 3.
    SEEA-R1 with training MGRM: Our proposed method, where the MGRM is trained (accuracy reported in Appendix) through the self-evolution process and then used to guide policy optimization. This configuration (e.g., green line) demonstrates consistent and significant performance improvement, approaching the results achieved with direct GT rewards.
The key finding from Figure 5 is that SEEA-R1, when utilizing its self-trained MGRM, can effectively learn and enhance its policy. This highlights the MGRM’s capability to provide reliable internal reward signals, enabling robust self-evolution and continuous performance gains without persistent dependence on external GT environment rewards.
Figure 5: Comparison of success rates of SEEA-R1 under different reward configurations.
@@ 4.2.4 Generalization on Out-of-Distribution Embodied Benchmark
Table 3: Success Rate in Embodiedeval benchmark
Model AttributeQA Navigation Object Interaction Social Interaction SpatialQA Overall
@@
Qwen2.5-VL-7B-Instruct* 34.74 18.97 5.62 0.00 21.15 18.29
@@
Qwen2.5-VL-7B-Instruct + MCTS + SFT 28.42 20.69 4.49 5.88 19.23 16.77
Qwen2.5-VL-7B-Instruct + MCTS + DPO 27.37 17.24 5.62 5.88 23.08 16.77
SEEA-R1 30.85 21.05 4.60 6.25 32.69 19.88
@@ 5 Conclusion
We introduced SEEA-R1, the first framework to adapt reinforcement fine-tuning (RFT) for training self-evolving embodied agents. To address the challenges of sparse rewards and limited generalization for embodied domains, SEEA-R1 integrates Tree-GRPO, which leverages MCTS to densify reward signals, and MGRM, a multi-modal reward model that generalizes across tasks and environments. SEEA-R1 achieves new state-of-the-art performance on ALFWorld, outperforming previous models, including GPT-4o, and demonstrates strong generalization in self-learning settings. These results demonstrate the potential of SEEA-R1 as a scalable framework for embodied reasoning and autonomous learning.
Limitation Despite its strong performance and generalization, SEEA-R1, like current embodied AI, does not yet fully address the complexities of operating in highly dynamic and unpredictable real-world environments. This reflects a broader, significant challenge for the field, indicating a substantial path forward for future research to bridge the gap to truly autonomous real-world agents. Future work might explore scaling experiments involving more diverse environments, larger sample size, and bigger models, aiming to enhance SEEA-R1’s capacity for dynamic environments.
6 Acknowledgment
@@ Appendix A Appendix Overview
  • Appendix G: Dataset Details
@@
  • Appendix H: Extended Experimental Results
@@
  • Appendix I: Case Studies on ALFWorld Tasks
    We present representative execution trajectories across various ALFWorld task categories (e.g., pick-and-place, look, clean, temperature manipulation) to highlight the SEEA-R1 agent’s capabilities. Visualizations of policy and reward model interactions are also included.
  • Appendix J: Prompts Used
@@ Appendix E Exploration of World Model as environmental feedback in Self-evolving Agent
Appendix F Code, Dataset and Extra Experiments
We include the detailed dataset construction and formulation in the supplement materials, together with more fine-grained ablation experiments on the model and data.
@@ Appendix G Dataset
Appendix H Experiments
H.1 Evaluation on the Genenal Multi-Modal Benchmark
@@
H.2 Impact of Sample Size and Batch Size on Training Stability
@@
H.3 Comparison with Different Algorithms
@@
Figure 10: Learning curve of Tree-GRPO over 30 training iterations
H.4 Long-Term Performance of Tree-GRPO
@@
H.5 Impact of Iterative Self-Evolution on MCTS Performance and Data Quality
@@
H.6 Evaluating Data Quality from Iterative MCTS + GRPO via Supervised Fine-Tuning
@@
Appendix I Case Studies on AlfWorld Tasks
@@
I.1 AlfWorld: Pick and Place Task
@@
I.2 AlfWorld: Multi-Object Placement Task
@@
I.3 AlfWorld: Look Task
@@
I.4 AlfWorld: Temperature Manipulation Tasks
Figures 13 and 14 illustrate representative instances of Heat tasks in the AlfWorld benchmark. In these scenarios, the agent is required to manipulate the thermal state of target objects—such as heating a mug prior to placing it in a coffee machine, or ensuring that a tomato is sufficiently warmed before disposal. Successfully completing these tasks demands the agent to not only identify the correct object and comprehend the temporal order of subgoals (e.g., "heat before use"), but also to correctly interact with relevant environmental appliances (e.g., microwaves or stoves). These tasks pose unique challenges in reasoning about object affordances, causal relationships, and multi-step transformations, as the agent must infer the implicit preconditions necessary for goal satisfaction and execute a valid sequence of actions accordingly.
I.5 AlfWorld: Clean Task
@@
I.6 AlfWorld: Composite Pick & Look Task
@@
I.7 Visualization of Policy and Reward Model Interactions
@@
Appendix J Prompts
J.1 Prompt for the ALFWorld Dataset

### v2

@@ 4.2.3 Comparative Study of Training Algorithms
4.2.4 Efficiency Analysis
Our method achieves higher success rates with lower computational cost (especially in inference token usage) via training-stage algorithmic optimization and lightweight inference design.
Search Algorithm Optimization.
Standard MCTS has complexity $O(T\cdot D\cdot K)$ (where $T$ = training steps, $D$ = search tree depth, $K$ = actions per node). We reduce this to $O(T\cdot max(D,K))$ via MCTS pruning as mentioned in Section 4.1.2.
Inference Efficiency.
[…]
Figure 5: Performance comparison of SEEA-R1 using different optimization algorithms on the multi-modal scenario of ALFWorld Benchmark over training iterations, more detailed figures are provided in Appendix G.3.
4.2.5 Ablation Study: Self-Evolution with Different Rewards
We conducted an ablation study to demonstrate that our SEEA-R1 can achieve sustained performance improvement by leveraging its internally trained Multi-Modal Generative Reward Model (MGRM), thereby reducing reliance on ground truth rewards from the simulator. We compared three critical configurations with results shown in Figure 6:
The key finding from Figure 6 is that SEEA-R1, when utilizing its self-trained MGRM (both in supervised and self-supervised paradigms), can effectively learn and enhance its policy. In the supervised paradigm (Figure 6(a)), the supervised MGRM enables steady performance that approaches the upper bound set by GT rewards. In the self-supervised paradigm, the self-supervised MGRM drives continuous success rate improvements, even outperforming the GT reward baseline in later iterations. This highlights the MGRM’s capability to provide reliable internal reward signals, enabling robust self-evolution and continuous performance gains without persistent dependence on external GT environment rewards.
Figure 6: Comparison of success rates of SEEA-R1 under different reward and training paradigms. (a) Analyzes the supervised paradigm, contrasting performance with GT reward, frozen MGRM, and supervised MGRM. (b) Examines the self-supervised paradigm, comparing GT reward, self-supervised MGRM, and the GPT-4o baseline.
@@ 4.2.6 Generalization on Out-of-Distribution Embodied Benchmark
Table 3: Success Rate in EmbodiedEval benchmark. Tasks include Attribute Question Answering (AttrQA), Spatial Question Answering (SpatialQA), Navigation, Object Interaction (Obj), and Social Interaction (Social).
Model AttrQA Navigation Obj Social SpatialQA Overall
@@
Vision-R1-7B zhan2025vision - - - - - 8.54
Ocean-R1 ming2025oceanr1 - - - - - 15.85
Qwen2.5-VL-7B* 34.74 18.97 5.62 0.00 21.15 18.29
@@
Qwen2.5-VL-7B + MCTS + SFT 28.42 20.69 4.49 5.88 19.23 16.77
Qwen2.5-VL-7B + MCTS + DPO 27.37 17.24 5.62 5.88 23.08 16.77
SEEA-R1 (w/ GT reward) 30.85 21.05 4.60 6.25 32.69 19.88
@@ 5 Conclusion
We introduced SEEA-R1, the first framework to adapt reinforcement fine-tuning (RFT) for training self-evolving embodied agents. To address the challenges of sparse rewards and limited generalization for embodied domains, SEEA-R1 integrates Tree-GRPO, which leverages MCTS to densify reward signals, and MGRM, a multi-modal reward model that generalizes across tasks and environments. SEEA-R1 achieves new state-of-the-art performance on ALFWorld, outperforming previous models, including GPT-4o, and demonstrates strong generalization in self-learning settings. These results demonstrate the potential of SEEA-R1 as a scalable framework for embodied reasoning and autonomous learning. Further discussions on the limitations of this work are in Appendix M.
6 Acknowledgments
@@ Appendix A Appendix Overview
  • Appendix F: Dataset Details
@@
  • Appendix G: Extended Experimental Results
@@
  • Appendix H: Case Studies on ALFWorld Tasks
    We present representative execution trajectories across various ALFWorld task categories (e.g., pick-and-place, look, clean, heat, cool) to highlight the SEEA-R1 agent’s capabilities. Visualizations of policy and reward model interactions are also included.
  • Appendix I: Qualitative Analysis of Reasoning Traces
    We analyze the evolution of reasoning capabilities by comparing the ReAct "Thought-Action-Observation" traces of the baseline (Qwen-2.5-VL-7B-Instruct) and SEEA-R1 (Iterations 5/10), focusing on the task "find two CDs and put them in the safe".
  • Appendix J: Detailed Real-World Tasks
    We detail all of tasks for each scenario in the real-world experiment.
  • Appendix K: Reward Model Accuracy Analysis
    We compare the task state judgment accuracy of the frozen MGRM (untrained Qwen-2.5-VL-7B-Instruct) and the GRPO-trained MGRM across "Success/Continue/Failure" states, with supporting data in Table 8.
  • Appendix L: Training Cost and Efficiency Analysis
    We specify the experimental hardware (8 NVIDIA A100 80GB GPUs) and frameworks (MS-Swift for training, vLLM for inference), quantify SEEA-R1’s time costs (sampling, policy training, total) under "with/without GT reward" configurations, and analyze efficiency trade-offs, with details in Table 10.
  • Appendix M: Limitations
    We discuss the limitations of SEEA-R1: like current embodied AI, it fails to fully address complexities in highly dynamic and unpredictable real-world environments (a major field-wide challenge). We also propose future directions, such as scaling experiments with more diverse environments, larger sample sizes, and bigger models to enhance its adaptability to dynamic settings.
  • Appendix N: Prompts Used
@@ Appendix D MCTS
D.1 Training and Efficiency Analysis
Training Cost.
[…]
Algorithmic Optimization.
Standard MCTS scales as $\mathcal{O}(b^{d})$ with branching factor $b$ and depth $d$. To improve scalability, SEEA-R1 introduces a pruning strategy combining probabilistic expansion and a path budget, effectively reducing complexity to $\mathcal{O}(pLk)$. Each node expands its $k$ actions with 50% probability, and full $k$-expansions are limited to $L=5$ per path.
Inference Efficiency.
[…]
Table 4: Inference Efficiency of SEEA-R1 on ALFWorld.
Model Success (%) Avg. Steps Total Tokens (Test)
Qwen2.5-VL-7B 11.57 28.10 10,701,255
SEEA-R1 36.19 23.37 6,554,090
@@ Appendix F Dataset Details
Appendix G Experiments
G.1 Evaluation on the Genenal Multi-Modal Benchmark
@@
G.2 Impact of Sample Size and Batch Size on Training Stability
@@
G.3 Comparison with Different Algorithms
@@
Figure 11: Learning curve of Tree-GRPO over 30 training iterations.
G.4 Long-Term Performance of Tree-GRPO
@@
G.5 Impact of Iterative Self-Evolution on MCTS Performance and Data Quality
@@
G.6 Evaluating Data Quality from Iterative MCTS + GRPO via Supervised Fine-Tuning
@@
Table 7: Performance comparison on the text-only scenario of ALFWorld Benchmark. ITO stands for Inference-Time Optimization. MPO (Meta Plan Optimization) xiong2025mpo provides the agent with the meta plan as context in inference time to improve performance.
Model w/o Training ITO ALFWorld Average
Seen Unseen
GPT-4o* achiam2023gpt - 78.6 83.6 81.1
Qwen2.5-7B* Yang2024Qwen25TR 71.4 75.4 73.4
Llama-3.1-8B* grattafiori2024llama - 22.9 28.4 25.7
Llama-3.1-70B* grattafiori2024llama - 78.6 73.9 76.3
GPT-4o* MPO 89.3 93.3 91.3
Llama-3.1-8B* MPO 50.0 52.2 51.1
Llama-3.1-8B + SFT* zeng2023agenttuning - 79.3 71.6 75.5
Llama-3.1-8B + ETO* song2024trial - 77.1 76.4 76.8
Llama-3.1-8B + KnowAgent* zhu2024knowagent - 80.0 74.9 77.5
Llama-3.1-8B + WKM* qiao2024agent - 77.1 78.2 77.7
Llama-3.1-8B + SFT* MPO 80.7 81.3 81.0
Llama-3.1-8B + ETO* MPO 85.0 79.1 82.1
SEEA-R1 (w/ GT-reward) - 85.3 85.1 85.2
@@
Appendix H Case Studies on ALFWorld Tasks
@@
H.1 ALFWorld: Pick and Place Task
@@
H.2 ALFWorld: Multi-Object Placement Task
@@
H.3 ALFWorld: Look Task
@@
Table 8: Accuracy Comparison Between Frozen and Training MGRM
Model Success Accuracy Continue Accuracy Failure Accuracy Overall Accuracy
Frozen MGRM 60.00% (9/15) 36.42% (460/1263) 81.25% (39/48) 51.59% (65/126)
Training MGRM 75.00% (9/12) 96.92% (1478/1525) 90.00% (27/30) 91.67% (77/84)
H.4 ALFWorld: Heat and Cool Tasks
Figures 14 and 15 illustrate representative instances of Heat tasks in the ALFWorld benchmark. In these scenarios, the agent is required to manipulate the thermal state of target objects—such as heating a mug prior to placing it in a coffee machine, or ensuring that a tomato is sufficiently warmed before disposal. Successfully completing these tasks demands the agent to not only identify the correct object and comprehend the temporal order of subgoals (e.g., "heat before use"), but also to correctly interact with relevant environmental appliances (e.g., microwaves or stoves). These tasks pose unique challenges in reasoning about object affordances, causal relationships, and multi-step transformations, as the agent must infer the implicit preconditions necessary for goal satisfaction and execute a valid sequence of actions accordingly.
H.5 ALFWorld: Clean Task
@@
H.6 ALFWorld: Composite Pick & Look Task
@@
H.7 Visualization of Policy and Reward Model Interactions
@@
Appendix I Qualitative Analysis of Reasoning Traces
This section provides qualitative examples of how the agent’s reasoning traces evolve throughout the training process, focusing on the task "find two CDs and put them in the safe" to illustrate improvements in reasoning capability.
I.0.1 Baseline Model (Qwen-2.5-VL-7B-Instruct)
The baseline model’s multi-round chain of thought exhibits limited adaptive reasoning, primarily relying on exhaustive search with repetitive failed actions:
Baseline (Qwen-2.5-VL-7B-Instruct) Reasoning Traces Thought: To find the CDs, I should start checking the shelves and drawers since they are likely places where CDs might be stored. I’$l$ begin with shelf 1. Action: go to shelf 1 Thought: There are no CDs on shelf 3 either. I should check the next shelf, which is shelf 4. Action: go to shelf 4 Observation: [image] Thought: Drawer 3 is closed. I need to open it to check inside. Action: open drawer 3 Observation: Nothing happens. … (more than ten consecutive failed attempts) Thought: Drawer 3 is still closed. I need to open it again to check inside. Action: open drawer 3 Observation: Nothing happens.
The baseline initially identifies plausible CD locations (shelves, drawers) and performs basic exploration but rigidly adheres to exhaustive search logic. It repeatedly executes unsuccessful actions (e.g., "open drawer 3" over ten times) without adjusting strategy, reflecting a lack of reflective reasoning about failure causes.
I.0.2 SEEA-R1 (Iteration 5)
After 5 training iterations, the model begins to exhibit adaptive reasoning, attempting alternative strategies when facing failures:
SEEA-R1 (Iteration 5) Reasoning Traces Thought: To find the CDs, I should start checking the shelves and drawers since they are likely places where CDs might be stored. Shelf 1 seems like a good place to start. Action: go to shelf 1 Thought: Safe 1 is open. I should put the CD inside. Action: put cd 1 in/on safe 1 Observation: Nothing happens. Thought: Putting the CD in the safe didn’t work. I should try closing the safe first. Action: close safe 1 Observation: Nothing happens. Thought: Closing the safe didn’t help either. I should try opening it again. Action: open safe 1 Observation: [image] Thought: Now that the safe is open, I should put the CD inside. Action: put cd 1 in/on safe 1
SEEA-R1 (Iteration 5) avoids mindless repetition of failed actions. When "put cd 1 in safe 1" fails, it reasons about potential causes (e.g., safe state) and tests alternatives (close → open). Though initial understanding of the safe’s state is flawed, it eventually succeeds through strategic adjustment, demonstrating emerging reflective reasoning.
I.0.3 SEEA-R1 (Iteration 10)
After 10 training iterations, the model exhibits structured planning and targeted corrective reasoning:
SEEA-R1 (Iteration 10) Reasoning Traces Thought: To solve the task, I need to find and take the first CD, then put it in the safe, then find and take the second CD, then put it in the safe. First, I need to find the first CD. A CD is more likely to appear in drawer, desk, shelf, and laundryhamper. I can check one by one, starting with drawer 1. Action: go to drawer 1 Action: put cd 1 in/on safe 1 Observation: Nothing happens. Thought: The action to put the first CD in the safe did not work, so I need to try opening the safe first. Action: open safe 1 Observation: [image] Action: put cd 1 in/on safe 1
SEEA-R1 (Iteration 10) shows significant improvements in planning and execution: (1) Its initial thought chain outlines a clear step-by-step plan (find → take → place CD1; repeat for CD2), with subsequent actions aligning closely with this plan. (2) When encountering failures (e.g., CD not placed), it immediately engages in reflective reasoning to identify fixes (e.g., "open the safe first"). (3) Unlike earlier stages, it minimizes redundant thoughts, focusing on actionable logic—evolving naturally toward efficient reasoning without explicit length penalties.
This progression confirms enhanced coherence and adaptability in the agent’s internal thought processes.
Appendix J Detailed Real-World Tasks
Table 9 lists all real-world tasks used in our experiments, organized by scenario. Each scenario contains 12 distinct tasks.
Home Living Room Children’s Room
1 Put the lemon on the sofa 1 Put the toy on the bedside table
2 Put the beverage on the TV cabinet 2 Put the toy on the bed
3 Put two fruits on the cabinet 3 Put the pillow on the bed
4 Put two fruits on the sofa 4 Close the bedside drawer
5 Put the beverage on the cabinet 5 Put the toy back on the shelf
6 Put the beverage on the sofa 6 Put the book on the bed
7 Put the apple in the plate 7 Put the book on the bookshelf
8 Throw the cup into the trash bin 8 Turn on the bedside lamp
9 Put the pear on the coffee table 9 Open the bedside drawer
10 Put the apple and pear on the coffee table 10 Put the book in the drawer
11 Put the apple in the refrigerator 11 Take the book out of the drawer
12 Tidy up items on the sofa 12 Put the toys on the bed into the drawer
Guest Room Kitchen
1 Put the book on the sofa 1 Put the pot on the gas stove
2 Put the book back on the bookshelf 2 Put the cherry tomatoes in the plate
3 Put the beverage on the sofa 3 Put the apple in the refrigerator
4 Throw the cup into the trash bin 4 Put the bowl in the sink
5 Put the apple in the plate 5 Wipe the countertop
6 Put the apple and banana on the coffee table 6 Put the empty plate in the cabinet
7 Throw the paper ball and cup into the trash bin 7 Put the sesame paste back into the refrigerator
8 Take the fruits on the sofa to the coffee table 8 Take the soybean paste to the stove
9 Put the apple and mango on the plate 9 Put the chili in the refrigerator
10 Put the pillow on the floor back on the sofa 10 Arrange the seasonings on the countertop neatly
11 Wipe the coffee table with a towel 11 Throw the trash on the countertop into the trash bin
12 Tidy up the pillows on the sofa 12 Put the tableware on the shelf
Tea Room Apartment
1 Put the book on the chair 1 Put the apple in the refrigerator
2 Put the book back on the bookshelf 2 Throw the beverage bottle into the trash bin
3 Put the beverage bottle on the bookshelf 3 Take the beverage from the refrigerator to the kitchen counter
4 Put the bowl on the table 4 Wipe the coffee table with a towel
5 Put the cushion on the chair 5 Take the beverage from the refrigerator to the table
6 Tidy up the tea sets on the tea table 6 Throw the paper ball and cup into the trash bin
7 Throw the trash on the table into the trash bin 7 Put the beverage and apple in the refrigerator
8 Put the water cup and beverage on the tea table 8 Put the clothes in the washing machine
9 Take a book and put it on the tea table 9 Put the apple and banana in the refrigerator
10 Wipe the tea table clean with a towel 10 Tidy up the pillows on the sofa
11 Put the cups on the tea table onto the dining table 11 Put the shoes on the shoe rack
12 Take the fruits on the bookshelf to the tea table 12 Put the apple in the refrigerator and the book on the sofa
Table 9: Detailed real-world tasks by scenario. Each scenario includes 12 tasks evaluating different manipulation capabilities.
Appendix K Reward Model Accuracy Analysis
The frozen reward model refers to directly using the untrained Qwen2.5-VL-7B-Instruct model to judge task states, and the accuracy of this untrained reward model in judging task states is only 51.59%. Therefore, the reward signals are highly unstable, as it can only provide correct reward signals in half of the cases. For example, when Embodied Agents are still in the process of completing the task, the actual task completion status is "Continue", but the reward model directly judges the task as "Failure", thus misjudging the actual situation and providing an incorrect reward signal.
In Table 8, the reward model trained with GRPO improves the prediction accuracy to 91.67%, enabling it to provide more accurate reward signals to Embodied Agents. The feedback from the frozen MGRM was primarily noisy rather than consistently rewarding incorrect actions, given its low accuracy of 51.59%—near random for binary judgments of task states (e.g., "Continue" vs. "Failure"). However, this noise included critical misjudgments that disrupted the agent’s learning process. Specifically, the frozen model frequently misclassified ongoing task states (where the agent should "Continue") as "Failure," prematurely signaling task termination. Over time, such erroneous feedback led the embodied agent to learn flawed behavioral patterns: it began to halt task execution prematurely, even when progress was valid, as it adapted to the unreliable reward signals. In short, while the frozen MGRM did not systematically reward incorrect actions, its noisy feedback—particularly misclassifications of ongoing progress as failure—introduced persistent biases that distorted the agent’s learned behaviors.
Appendix L Training Cost and Efficiency Analysis
This section provides detailed analysis of the training costs and efficiency of our proposed methods, including experimental setup, computational requirements, and performance comparisons.
L.1 Experimental Environment
All experiments were conducted on a high-performance computing cluster equipped with 8 NVIDIA A100 80GB GPUs. We used the MS-Swift framework for distributed model training, which provided efficient scaling across multiple GPUs. For inference performance evaluation, we employed the vLLM library kwon2023efficient to ensure high-throughput and low-latency model serving.
L.2 Training Efficiency Comparison
Table 10 presents a comprehensive comparison of the training efficiency between different configurations of our SEEA-R1 model. All time metrics in the table represent average values per training iteration; specifically, they include average sampling time per iteration, policy training time per iteration, reward model training time per iteration, and total training time per iteration. All time units are standardized to minutes for consistency.
Table 10: Training Efficiency Comparison of SEEA-R1 Under Different Reward Configurations (Per Training Iteration).
Method MGRM Training Method Average Sampling Time Policy Training Time Reward Model Training Time Total Time
SEEA-R1 (w/o GT-reward) TTRL 173.5 mins 4.0 mins 244.0 mins 421.5 mins
SEEA-R1 (w/ GT-reward) - 190.1 mins 4.0 mins - 194.1 mins
L.3 Key Observations
From the training efficiency analysis, we can draw several important conclusions:
1. Self-supervised Training Overhead: The SEEA-R1 model with self-supervised MGRM training (w/o GT-reward) requires significantly more total training time (248 minutes) compared to the supervised configuration, primarily due to the additional reward model training time (244 minutes).
2. Sampling Efficiency: Despite the longer total training time, the self-supervised configuration demonstrates more efficient sampling, with an average sampling time of 173.50 minutes compared to 190.12 minutes for the supervised configuration.
3. Policy Training Consistency: The policy training time remains consistent at 4 minutes across both configurations, indicating that the core policy learning process is not significantly affected by the reward configuration.
4. Cost-Effectiveness Trade-off: The self-supervised approach, while computationally more expensive, offers the advantage of not requiring ground truth rewards, making it more practical for real-world scenarios where labeled data is scarce or unavailable.
These findings highlight the computational considerations involved in deploying our proposed methods and provide insights into the trade-offs between training efficiency and the need for ground truth supervision.
Appendix M Limitations
Despite its strong performance and generalization, SEEA-R1, like current embodied AI, does not yet fully address the complexities of operating in highly dynamic and unpredictable real-world environments. This reflects a broader, significant challenge for the field, indicating a substantial path forward for future research to bridge the gap to truly autonomous real-world agents. Future work might explore scaling experiments involving more diverse environments, larger sample size, and bigger models, aiming to enhance SEEA-R1’s capacity for dynamic environments.
Appendix N Prompts
N.1 Prompt for the ALFWorld Dataset